# Quantum metrics from length functions on quantum groups

### v1

@@ 4. The conjugation coaction
We have thus shown that $E$ maps $L^{\infty}({\mathbb{G}})$ to itself, and by Lemma 4.1 it therefore takes values in $L^{\infty}_{z}({\mathbb{G}})$. The restriction of $E$ is thus the unique $h$-preserving conditional expectation from $L^{\infty}({\mathbb{G}})$ onto $L^{\infty}_{z}({\mathbb{G}})$. By [AC17, Theorem 3.7] we have $L^{\infty}_{z}({\mathbb{G}})=\operatorname{Pol}_{z}({\mathbb{G}})^{\prime\prime}$ and $E$ is therefore identical to the conditional expectation considered in the proof of [Wan17, Lemma 6.3]. In particular, we have the formula
@@
derived in the proof of [Wan17, Lemma 6.3], and hence that $E$ maps $\operatorname{Pol}({\mathbb{G}})$ to $\operatorname{Pol}_{z}({\mathbb{G}})$ and, by continuity, $C({\mathbb{G}})$ to $C_{z}({\mathbb{G}})$. For the convenience of the reader, we now derive the formula (4.3) directly. To this end, first note that since $\Lambda(1)$ is separating for $L^{\infty}({\mathbb{G}})$ it suffices to show that
@@
for all $\beta\in\operatorname{Irred}({\mathbb{G}})$ and $p,q\in\{1,\dots,d_{\beta}\}$. Note also that since ${\mathbb{G}}$ is assumed Kac, the orthogonality relations [Tim08, Proposition 5.3.8 (iii)] simplify to the expression
$h(u_{ij}^{\alpha}(u_{pq}^{\beta})^{*})=\delta_{\alpha,\beta}\delta_{ip}\delta_{jq}d_{\alpha}^{-1}.$
@@
$=\delta_{\alpha,\beta}\delta_{i,j}\delta_{pq}d_{\alpha}^{-2}$
@@ 4.1. The conditional expectation is a slip-norm contraction
Our next aim is to show that the conditional expectation introduced in the previous section is a contraction for the seminorm $L_{\ell}$ associated with a proper length function $\ell$. Let therefore ${\mathbb{G}}$ be a compact quantum group of Kac type, and fix a proper length function $\ell\colon\operatorname{Irred}({\mathbb{G}})\to[0,\infty)$. On $L^{2}({\mathbb{G}})\hat{\otimes}L^{2}({\mathbb{G}})$ we consider the densely defined symmetric operator
${\operatorname{id}}\otimes\mathcal{D}_{\ell}\colon\Lambda(\operatorname{Pol}({\mathbb{G}}))\odot\Lambda(\operatorname{Pol}({\mathbb{G}}))$ $\longrightarrow L^{2}({\mathbb{G}})\hat{\otimes}L^{2}({\mathbb{G}})$
@@
$\left\langle\partial_{\ell}(E(a))\Lambda(x),\Lambda(y)\right\rangle=$
$=\left\langle\mathcal{D}_{\ell}((h\otimes{\operatorname{id}})\delta(a))\Lambda(x),\Lambda(y))\right\rangle-\left\langle((h\otimes{\operatorname{id}})\delta(a))\mathcal{D}_{\ell}\Lambda(x),\Lambda(y))\right\rangle$
$=\left\langle\delta(a)(\Lambda(1)\otimes\Lambda(x)),\Lambda(1)\otimes\mathcal{D_{\ell}}\Lambda(y)\right\rangle-\left\langle\delta(a)(\Lambda(1)\otimes\mathcal{D}_{\ell}\Lambda(x)),\Lambda(1)\otimes\Lambda(y)\right\rangle$
$=\left\langle Z^{*}(1\otimes a)Z(\Lambda(1)\otimes\Lambda(x)),\Lambda(1)\otimes\mathcal{D_{\ell}}\Lambda(y)\right\rangle-\left\langle Z^{*}(1\otimes a)Z(\Lambda(1)\otimes\mathcal{D}_{\ell}\Lambda(x)),\Lambda(1)\otimes\Lambda(y)\right\rangle$
$=\left\langle Z^{*}(1\otimes\mathcal{D_{\ell}}a)Z(\Lambda(1)\otimes\Lambda(x)),\Lambda(1)\otimes\Lambda(y)\right\rangle-\left\langle Z^{*}(1\otimes a\mathcal{D}_{\ell})Z(\Lambda(1)\otimes\Lambda(x)),\Lambda(1)\otimes\Lambda(y)\right\rangle$
$=\left\langle Z^{*}(1\otimes\partial_{\ell}(a)Z(\Lambda(1)\otimes\Lambda(x)),\Lambda(1)\otimes\Lambda(y))\right\rangle$
$=\left\langle(h\otimes{\operatorname{id}})\delta(\partial_{\ell}(a))\Lambda(x),\Lambda(y)\right\rangle$
$=\left\langle E(\partial_{\ell}(a))\Lambda(x),\Lambda(y)\right\rangle.$
@@
The last bit of information about $L_{\ell}$ needed for the proof of Theorem A, is that it satisfies Li’s left-invariance condition [Li09]; i.e. that the following result holds true.
@@
The bounded extension of $[\mathcal{D}_{\ell},(\varphi\otimes{\operatorname{id}})\Delta(a)]$, i.e. the operator $\partial_{\ell}((\varphi\otimes{\operatorname{id}})\Delta(a))$, therefore agrees with $(\varphi\otimes{\operatorname{id}})(W^{*}(1\otimes\partial_{\ell}(a))W)$, from which the claim now follows via the following estimates:
@@ 5. Central approximations of the counit
When ${\mathbb{G}}$ is of Kac type then $\varphi\in{\mathcal{S}}(C({\mathbb{G}}))$ is central if and only if $\varphi=\varphi\circ E$, where $E$ is the conditional expectation $E\colon C({\mathbb{G}})\to C_{z}({\mathbb{G}})$ from Proposition 4.2.
@@
$\varphi\circ E(u_{ij}^{\alpha})=\varphi\big{(}\delta_{i,j}d_{\alpha}^{-1}\chi(u^{\alpha})\big{)}=\delta_{i,j}d_{\alpha}^{-1}\sum_{k=1}^{d_{\alpha}}\varphi(u_{kk}^{\alpha})=\delta_{i,j}\varphi(u_{ii})=\varphi(u_{ij}^{\alpha}).$
@@
This result can be derived in numerous ways from the existing characterisations of coamenability (see e.g.  [Tom06]). Here the result will be derived using Følner sequences, since this approach makes the connection to the corresponding result for discrete groups quite transparent.
@@
Denote by $F_{n}\subset\operatorname{Irred}({\mathbb{G}})$ a Følner sequence for ${\mathbb{G}}$ [Kye08], and define a sequence of functions $\omega_{n}\colon\operatorname{Irred}({\mathbb{G}})\to[0,\infty)$ by
@@
Then [HWW24, Lemma 5.5] shows that $\lim_{n}\omega_{n}(\gamma)=1$ for all $\gamma\in\operatorname{Irred}({\mathbb{G}})$ and that the associated multipliers $T_{n}\colon\operatorname{Pol}({\mathbb{G}})\to\operatorname{Pol}({\mathbb{G}})$
@@
extend to ucp maps on $L^{\infty}({\mathbb{G}})$, and hence also on $C({\mathbb{G}})$ since $T_{n}$ preserves $\operatorname{Pol}({\mathbb{G}})$. Setting $\chi_{n}:=\epsilon\circ T_{n}\colon C({\mathbb{G}})\to{\mathbb{C}}$ we therefore obtain a sequence of states satisfying
@@
$\chi_{n}(u_{ij}^{\alpha})=\omega_{n}(\alpha)\delta_{ij}\underset{n\to\infty}{\longrightarrow}\delta_{ij}=\epsilon(u_{ij}^{\alpha}).$
[…]
@@
Let ${\mathbb{G}}$ be a compact, coamenable quantum group of Kac type and let $\ell\colon\operatorname{Irred}({\mathbb{G}})\to[0,\infty)$ be a proper length function. For any two central states, $\varphi$ and $\psi$, on $C({\mathbb{G}})$ it holds that
@@
$=:d_{\ell}^{z}\big{(}\varphi\restriction_{C_{z}({\mathbb{G}})},\psi\restriction_{C_{z}({\mathbb{G}})}\big{)}$
@@
For a compact quantum group ${\mathbb{G}}$ and $\chi\in{\mathcal{S}}(C({\mathbb{G}}))$ we denote by $\beta_{\chi}\colon C({\mathbb{G}})\to C({\mathbb{G}})$ the ucp map defined by $\beta_{\chi}(a)=({\operatorname{id}}\otimes\chi)\Delta(a)$.
@@
Let ${\mathbb{G}}$ be a compact, coamenable quantum group and let $\ell\colon\operatorname{Irred}({\mathbb{G}})\to[0,\infty)$ be a proper length function. For any $\chi\in{\mathcal{S}}(C({\mathbb{G}}))$ and any $a\in C({\mathbb{G}})$ it holds that $\|x-\beta_{\chi}(a)\|\leqslant d_{\ell}(\epsilon,\chi)\cdot L_{\ell}(a)$.
We remark that the estimate in Lemma 5.5, as well as its proof, is heavily inspired by the analysis of the quantum metric structure of $q$-deformations, studied in [KK25, AKK22a, AKK22b].
@@
We aim to apply Kaad’s characterisation of compact quantum metric spaces presented in Theorem 2.8. To meet the criteria set forth in Theorem 2.8, it suffices to show that $(C({\mathbb{G}}),L_{\ell})$ has finite diameter (in the sense of Definition 2.7) and provide a sequence of finite rank ucp maps $\beta_{n}\colon C({\mathbb{G}})\to C({\mathbb{G}})$, which converges to the identity uniformly on the $L_{\ell}$-unit ball. To this end, consider again the sequence of central states, $(\chi_{n})_{n\in{\mathbb{N}}}$ provided by Lemma 5.2. We augment the sequence by setting $\chi_{0}:=h$ and put $\beta_{n}:=\beta_{\chi_{n}}$; see Definition 5.4. Using the the centrality of $\chi_{n}$ and $\epsilon$ together with Lemma 5.5 and Lemma 5.3 we the obtain:
$\|a-\beta_{n}(a)\|\leqslant d_{\ell}(\epsilon,\chi_{n})\cdot L(a)=d_{\ell}^{z}(\epsilon\restriction_{C_{z}({\mathbb{G}})},\chi_{n}\restriction_{C_{z}({\mathbb{G}})})\cdot L(a),$ (5.2)
@@
$d_{\ell}^{z}(\epsilon\restriction_{C_{z}({\mathbb{G}})},\chi_{n}\restriction_{C_{z}({\mathbb{G}})})<\infty\textup{ for all $n\in{\mathbb{N}}_{0}$ and }\lim_{n\to\infty}d_{\ell}^{z}(\epsilon\restriction_{C_{z}({\mathbb{G}})},\chi_{n}\restriction_{C_{z}({\mathbb{G}})})=0.$
@@
$\|x-h(x)1\|=\|x-\beta_{0}(x)\|\leqslant d_{\ell}^{z}(\epsilon\restriction_{C_{z}({\mathbb{G}})},h\restriction_{C_{z}({\mathbb{G}})}))\cdot L(x),$
and hence $(C({\mathbb{G}}),L_{\ell})$ has finite diameter. Moreover, since $\beta_{n}$ is ucp with finite-dimensional image and $\lim_{n\to\infty}d_{\ell}^{z}(\epsilon\restriction_{C_{z}({\mathbb{G}})},\chi_{n}\restriction_{C_{z}({\mathbb{G}})})=0$, the conditions in [Kaa24, Theorem 3.1] are fulfilled and the proof is complete. ∎
@@
on which the formula defining $L_{\ell}$ yields a finite number. It is possible that the techniques used to prove Theorem A and Corollary B can be modfied to show the corresponding (and stronger) results for the maximal version of $L_{\ell}$, but to minimise the technicalities stemming from unbounded operator theory we shall not pursue this question further.

### v3

@@ 4. The conjugation coaction
We have thus shown that $E$ maps $L^{\infty}({\mathbb{G}})$ to itself, and by Lemma 4.1 it therefore takes values in $L^{\infty}_{z}({\mathbb{G}})$. The restriction of $E$ is thus the unique $h$-preserving conditional expectation from $L^{\infty}({\mathbb{G}})$ onto $L^{\infty}_{z}({\mathbb{G}})$. By [Crann:CaracterDensity, Theorem 3.7] we have $L^{\infty}_{z}({\mathbb{G}})=\operatorname{Pol}_{z}({\mathbb{G}})^{\prime\prime}$ and $E$ is therefore identical to the conditional expectation considered in the proof of [Wang:lacunary, Lemma 6.3]. In particular, we have the formula
@@
derived in the proof of [Wang:lacunary, Lemma 6.3], and hence that $E$ maps $\operatorname{Pol}({\mathbb{G}})$ to $\operatorname{Pol}_{z}({\mathbb{G}})$ and thus, by continuity, $C({\mathbb{G}})$ to $C_{z}({\mathbb{G}})$. For the convenience of the reader, we now derive the formula (4.3) directly. To this end, first note that since $\Lambda(1)$ is separating for $L^{\infty}({\mathbb{G}})$ it suffices to show that
@@
for all $\beta\in\operatorname{Irred}({\mathbb{G}})$ and $p,q\in\{1,\dots,d_{\beta}\}$. Note also that since $\mathbb{G}$} is assumed Kac, the orthogonality relations [Timmermann-book, Proposition 5.3.8 (iii)] simplify to the expression
$h((u_{ij}^{\alpha})^{*}u_{pq}^{\beta})=\delta_{\alpha,\beta}\delta_{ip}\delta_{jq}d_{\alpha}^{-1}.$
@@
$=\delta_{\alpha,\beta}\delta_{i,j}\delta_{p,q}d_{\alpha}^{-2}$
@@ 4.1. The conditional expectation is a slip-norm contraction
Our next aim is to show that the conditional expectation introduced in the previous section is a contraction for the seminorm $L_{\ell}$ associated with a proper length function $\ell$. Let therefore $\mathbb{G}$} be a compact quantum group of Kac type, and fix a proper length function $\ell\colon\operatorname{Irred}({\mathbb{G}})\rightarrow[0,\infty)$. On $L^{2}({\mathbb{G}})\hat{\otimes}L^{2}({\mathbb{G}})$ we consider the densely defined symmetric operator
${\operatorname{id}}\otimes\mathcal{D}_{\ell}\colon\Lambda(\operatorname{Pol}({\mathbb{G}}))\odot\Lambda(\operatorname{Pol}({\mathbb{G}}))$ -$\mathrel{\smash{\meno}}\mathrel{\mkern-3.0mu}\rightarrow L^{2}({\mathbb{G}})\hat{\otimes}L^{2}({\mathbb{G}})$
@@
$\left\langle\Lambda(x),\partial_{\ell}(E(a))\Lambda(y)\right\rangle=$
$=\left\langle\Lambda(x),\mathcal{D}_{\ell}((h\otimes{\operatorname{id}})\delta(a))\Lambda(y))\right\rangle-\left\langle\Lambda(x),((h\otimes{\operatorname{id}})\delta(a))\mathcal{D}_{\ell}\Lambda(y))\right\rangle$
$=\left\langle\Lambda(1)\otimes\mathcal{D_{\ell}}\Lambda(x),\delta(a)(\Lambda(1)\otimes\Lambda(y))\right\rangle-\left\langle\Lambda(1)\otimes\Lambda(x),\delta(a)(\Lambda(1)\otimes\mathcal{D}_{\ell}\Lambda(y))\right\rangle$
$=\left\langle\Lambda(1)\otimes\mathcal{D_{\ell}}\Lambda(x),Z^{*}(1\otimes a)Z(\Lambda(1)\otimes\Lambda(y))\right\rangle-\left\langle\Lambda(1)\otimes\Lambda(x),Z^{*}(1\otimes a)Z(\Lambda(1)\otimes\mathcal{D}_{\ell}\Lambda(y))\right\rangle$
$=\left\langle\Lambda(1)\otimes\Lambda(x),Z^{*}(1\otimes\mathcal{D_{\ell}}a)Z(\Lambda(1)\otimes\Lambda(y))\right\rangle-\left\langle\Lambda(1)\otimes\Lambda(x),Z^{*}(1\otimes a\mathcal{D}_{\ell})Z(\Lambda(1)\otimes\Lambda(y))\right\rangle$
$=\left\langle\Lambda(1)\otimes\Lambda(x),Z^{*}(1\otimes\partial_{\ell}(a)Z(\Lambda(1)\otimes\Lambda(y))\right\rangle$
$=\left\langle\Lambda(x),(h\otimes{\operatorname{id}})\delta(\partial_{\ell}(a))\Lambda(y)\right\rangle$
$=\left\langle\Lambda(x),E(\partial_{\ell}(a))\Lambda(y)\right\rangle.$
@@
The last bit of information about $L_{\ell}$ needed for the proof of Theorem A, is that it satisfies Li’s left-invariance condition [Li:ECQ]; i.e. that the following result holds true.
@@
The bounded extension of $\mathcal{D}_{\ell},(\varphi\otimes{\operatorname{id}})\Delta(a)]$, i.e. the operator $\partial_{\ell}((\varphi\otimes{\operatorname{id}})\Delta(a))$, therefore agrees with $(\varphi\otimes{\operatorname{id}})(W^{*}(1\otimes\partial_{\ell}(a))W)$, from which the claim now follows via the following estimates:
@@ 5. Central approximations of the counit
When $\mathbb{G}$} is of Kac type then $\varphi\in{\mathcal{S}}(C({\mathbb{G}}))$ is central if and only if $\varphi=\varphi\circ E$, where $E\colon C({\mathbb{G}})\rightarrow C_{z}({\mathbb{G}})$ is the conditional expectation from Proposition 4.2.
@@
$\varphi\circ E(u_{ij}^{\alpha})=\varphi\big(\delta_{i,j}d_{\alpha}^{-1}\chi(u^{\alpha})\big)=\delta_{i,j}d_{\alpha}^{-1}\sum_{k=1}^{d_{\alpha}}\varphi(u_{kk}^{\alpha})=\delta_{i,j}\varphi(u_{ii}^{\alpha})=\varphi(u_{ij}^{\alpha}).$
@@
This result can be derived in numerous ways from the existing characterisations of coamenability (see e.g.  [tomatsu-amenable]). Here the result will be derived using Følner sequences, since this approach makes the connection to the corresponding result for discrete groups transparent.
@@
Denote by $F_{n}\subset\operatorname{Irred}({\mathbb{G}})$ a Følner sequence for $\mathbb{G}$} [DK:L2-of-coamenable], and define a sequence of functions $\omega_{n}\colon\operatorname{Irred}({\mathbb{G}})\rightarrow[0,\infty)$ by
@@
Then [HWW:Pointwise-convergence, Lemma 5.5] shows that $\lim_{n}\omega_{n}(\gamma)=1$ for all $\gamma\in\operatorname{Irred}({\mathbb{G}})$ and that the associated multipliers $T_{n}\colon\operatorname{Pol}({\mathbb{G}})\rightarrow\operatorname{Pol}({\mathbb{G}})$
@@
extend to ucp maps on $L^{\infty}({\mathbb{G}})$, and hence also on $C({\mathbb{G}})$ since $T_{n}$ preserves $\operatorname{Pol}({\mathbb{G}})$. Setting $\chi_{n}:=\epsilon\circ T_{n}\colon C({\mathbb{G}})\rightarrow{\mathbb{C}}$ we therefore obtain a sequence of states satisfying
@@
$\chi_{n}(u_{ij}^{\alpha})=\omega_{n}(\alpha)\delta_{ij}\underset{n\rightarrow\infty}{\mathrel{\smash{\meno}}\mathrel{\mkern-3.0mu}\rightarrow}\delta_{ij}=\epsilon(u_{ij}^{\alpha}).$
By linearity, the sequence of states $\chi_{n}$ therefore converges to $\epsilon$ pointwise on $\operatorname{Pol}({\mathbb{G}})$ and by density also on $C({\mathbb{G}})$. Lastly, to see that $\chi_{n}$ is only supported in finitely many matrix coefficients, note that $\chi_{n}(u_{ij}^{\gamma})\neq 0$ if and only if $N_{\alpha,\bar{\beta}}^{\gamma}\neq 0$ for some $\alpha,\beta\in F_{n}$. But since $F_{n}$ is finite, the union of supports of products of the form $\alpha \cdot\bar{\beta}$ with $\alpha,\beta\in F_{n}$ is also finite, so when $\gamma$ falls outside this finite set we have $\chi_{n}(u_{ij}^{\gamma})=0$. ∎
The main reason why the central approximation $\chi_{n}$ of the counit provided by Lemma 5.2 is relevant to us, is contained in the following lemma, which shows that the distance between $\chi_{n}$ and counit can be measured by their restrictions to the algebra of central functions.
@@
Let $\mathbb{G}$} be a compact, coamenable quantum group of Kac type and let $\ell\colon\operatorname{Irred}({\mathbb{G}})\rightarrow[0,\infty)$ be a proper length function. For any two central states, $\varphi$ and $\psi$, on $C({\mathbb{G}})$ it holds that
@@
$=:d_{\ell}^{z}\big(\varphi\hskip-4.26773pt\restriction_{C_{z}({\mathbb{G}})},\psi\hskip-4.26773pt\restriction_{C_{z}({\mathbb{G}})}\big)$
@@
For a compact quantum group $\mathbb{G}$} and $\chi\in{\mathcal{S}}(C({\mathbb{G}}))$ we denote by $\beta_{\chi}\colon C({\mathbb{G}})\rightarrow C({\mathbb{G}})$ the ucp map defined by $\beta_{\chi}(a)=({\operatorname{id}}\otimes\chi)\Delta(a)$.
@@
Let $\mathbb{G}$} be a compact, coamenable quantum group and let $\ell\colon\operatorname{Irred}({\mathbb{G}})\rightarrow[0,\infty)$ be a proper length function. For any $\chi\in{\mathcal{S}}(C({\mathbb{G}}))$ and any $a\in C({\mathbb{G}})$ it holds that $\|a-\beta_{\chi}(a)\|\leqslant d_{\ell}(\epsilon,\chi)\cdot L_{\ell}(a)$.
We remark that the estimate in Lemma 5.5, as well as its proof, is heavily inspired by the analysis of the quantum metric structure of $q$-deformations, studied in [KaadKyed2022, AKK:Podcon, AKK:PolyApprox].
@@
We aim to apply Kaad’s characterisation of compact quantum metric spaces presented in Theorem 2.8. To meet the criteria set forth in Theorem 2.8, it suffices to show that $(C({\mathbb{G}}),L_{\ell})$ has finite diameter (in the sense of Definition 2.7) and provide a sequence of finite rank ucp maps $\beta_{n}\colon C({\mathbb{G}})\rightarrow C({\mathbb{G}})$, which converges to the identity uniformly on the $L_{\ell}$-unit ball. To this end, consider again the sequence of central states, $(\chi_{n})_{n\in{\mathbb{N}}}$ provided by Lemma 5.2. We augment the sequence by setting $\chi_{0}:=h$ and put $\beta_{n}:=\beta_{\chi_{n}}$; see Definition 5.4. Using the centrality of $\chi_{n}$ and $\epsilon$ together with Lemma 5.5 and Lemma 5.3 we the obtain:
$\|a-\beta_{n}(a)\|\leqslant d_{\ell}(\epsilon,\chi_{n})\cdot L_{\ell}(a)=d_{\ell}^{z}(\epsilon\hskip-4.26773pt\restriction_{C_{z}({\mathbb{G}})},\chi_{n}\hskip-4.26773pt\restriction_{C_{z}({\mathbb{G}})})\cdot L_{\ell}(a),$ (5.2)
@@
$d_{\ell}^{z}(\epsilon\hskip-4.26773pt\restriction_{C_{z}({\mathbb{G}})},\chi_{n}\hskip-4.26773pt\restriction_{C_{z}({\mathbb{G}})})<\infty\textup{ for all $n\in{\mathbb{N}}_{0}$ and }\lim_{n\rightarrow\infty}d_{\ell}^{z}(\epsilon\hskip-4.26773pt\restriction_{C_{z}({\mathbb{G}})},\chi_{n}\hskip-4.26773pt\restriction_{C_{z}({\mathbb{G}})})=0.$
@@
$\|a-h(a)1\|=\|a-\beta_{0}(a)\|\leqslant d_{\ell}^{z}(\epsilon\hskip-4.26773pt\restriction_{C_{z}({\mathbb{G}})},h\hskip-4.26773pt\restriction_{C_{z}({\mathbb{G}})}))\cdot L_{\ell}(a)$},
and hence $(C({\mathbb{G}}),L_{\ell})$ has finite diameter. Moreover, since $\beta_{n}$ is ucp with finite-dimensional image and $\lim_{n\rightarrow\infty}d_{\ell}^{z}(\epsilon\hskip-4.26773pt\restriction_{C_{z}({\mathbb{G}})},\chi_{n}\hskip-4.26773pt\restriction_{C_{z}({\mathbb{G}})})=0$, the conditions in [Kaad2023, Theorem 3.1] are fulfilled and the proof is complete. ∎
@@
on which the formula defining $L_{\ell}$ yields a finite number. It is possible that the techniques used to prove Theorem A and Corollary B can be modified to show the corresponding (and stronger) results for the maximal version of $L_{\ell}$, but to minimise the technicalities stemming from unbounded operator theory, we shall not pursue this question further.